OpenStudy (anonymous):

The first term of an arithmetic sequence is -3 and the fifteenth term is 53. What is the common difference of the sequence? 14/13 25/7 4

OpenStudy (anonymous):

A

OpenStudy (cwrw238):

53 - (-3) = 56 and the common difference has been added 14 times to the first term CD = 56 / 14 = 4